Question 5 (1 point)

Experimenters injected a growth hormone gene into thousands of carp eggs. Of the 264 carp that grew from these eggs, 30 incorporated the gene into their DNA (Science News, May 20, 1989). With a confidence of 90%, what is the margin of error for the proportion of all carp that would incorporate the gene into their DNA?

Question 5 options:

1)

0.0006

2)

0.0195

3)

0.0320

4)

0.0020

5)

0.0250
Answer #1

solution:- option 3)

explanation:-

given that x = 30 and n = 264

proportion p^ = x/n = 30/264

= 0.114

the value of 90% confidence from z table is 1.645

margin of error formula

=> z * sqrt(p*(1-p)/n)

=> 1.645 * sqrt(0.114*(1-0.114)/264)

=> 0.0320